# Spatially resolved broad line region in a quasar at z=4

### v1

@@ 5.1 BLR geometry and size
The fitted parameters summarised in Table 4 show that % of the clouds are in the outflow, and that this component therefore completely dominates the BLR. The remaining 17% of the clouds are in the disky component on bound elliptical orbits. Our line of sight is near the top edge of the outflow, because the inclination is from face-on while the total angle from the mid-plane subtended by BLR clouds is . We note that if the line of sight was oriented further into the outflowing wedge, one might expect to see broad absorption lines in the rest-frame optical spectrum (Elvis, 2000) but these are not observed (Wolf et al., 2024). The emission is anisotropic with indicating that the side of the clouds illuminated by the AGN emits preferentially. Equally important is the mid-plane opacity with meaning that the emission from clouds behind the mid-plane is almost entirely blocked, consistent both with the basis of the FRADO model in which a dusty disk extends to small radii at low scale heights and also the optically thick nature of super-Eddington accretion disks. Together these lead to the configuration in Fig. 6, where the clouds are coloured according to their line-of-sight velocity and have a size corresponding to their apparent brightness.
@@
The average radius for the BLR is pc for H and pc for H. While these are the same within their uncertainties (see Table 4 and Fig. 9), this difference is expected both observationally, based on RM delays for different lines (Bentz et al., 2010), and on modelling of radiation pressure confined BLR clouds (Netzer, 2020).
@@
While the size appears to match that inferred by applying a correction to the R-L relation favoured by Du & Wang (2019), this does not provide an explanation of why is so low. Indeed, from the plot of FWHM versus in Shen & Ho (2014), both of these values for J0529 seem rather typical of QSOs generally. Instead we propose that much of the deviation from the scaling relations that has previously been attributed to luminosity or accretion rate, could, analogous to the C IV line, be due to the impact of the outflow on the H I line properties. Indeed, the disky part of the BLR in Fig. 5 contains only 27% of the total line luminosity. Applying this to the typical equation which has a typical form where is usually taken to represent the total AGN luminosity. If instead we consider only the luminosity associated with the disky part of the BLR tracing rotationally supported gas (using the line luminosity as a suitable proxy), then the inferred size is reduced by a factor . This would bring the estimate from Bentz et al. (2013) into good agreement with our measurement.
@@
Similarly, using only the disky part of the BLR would bring the equivalent width down to Å, within the range of typical values. And we would find which is what one would expect for a highly accreting QSO, particularly in the context of Eddington ratio discussed below. However, applying this new value would lead to a revised BLR size from Du & Wang (2019) that is much too small. We therefore speculate that the over-estimate of the BLR size that has been discussed in the literature for highly accreting and highly luminous QSOs could be related to the impact of BLR outflows on the line properties.
@@ 5.2 Dynamical black hole mass
The form of the scaling relation used to estimate black hole mass is typically where represents AGN luminosity and (or FWHM) the line width. If, as above, we take to be a proxy for the line luminosity (noting that line luminosity is used explicitly in some relations), then applying this to only the disky part of the BLR profile which has km -1 and contains only 27% of the luminosity, would reduce the inferred by a factor 14 (or a factor 5 if using FWHM, due to the differing FWHM/ ratios). This use of the rotational velocity of the BLR rather than the full line width (which is broadened by the outflow) would bring all the single epoch estimates much closer to the BH mass we have measured, and supports the speculation above that the outflow could be a major contributor to the bias in single epoch estimates for luminous QSOs
@@ 5.3 Eddington ratio
The Eddington ratio is determined by the black-hole mass and the bolometric luminosity , which is commonly derived from a monochromatic luminosity and a bolometric correction . Also, the emission from the accretion disc requires an anisotropy correction that depends on the viewing angle, and to a smaller extent on the black-hole spin since the deflection of light due to gravitational lensing is stronger for disc annuli that are closer to the black hole (Nemmen & Brotherton, 2010; Lai et al., 2023). Generic values based on a single average spectral energy distribution (SED; Richards et al. 2006) are usually appropriate when the disc is large compared to the innermost stable circular orbit (ISCO) around the black hole. While the overall size and surface area of a disc is set by its luminosity, black holes of increasing mass will have the hole inside the ISCO grow proportionally to mass. Discs around higher-mass black holes thus lack the hottest and most UV-emitting parts, which are otherwise present in discs of similar luminosity around lower-mass black holes. This implies a non-linear dependence of on mass and disc luminosity, which was explored in the framework of thin-disc models by several authors (e.g. Nemmen & Brotherton, 2010; Netzer, 2019; Lai et al., 2023). These explorations tend to offer a mean relation for as a function of monochromatic luminosity assuming a mean related black-hole mass or Eddington ratio.
@@
In our case of J0529, we use the 5100 Å luminosity from Sec. 3 and the 33∘ inclination implied by our model. The mean relations by the three aforementioned works result in values ranging from 2.2 to 2.7, but are not applicable in the case of J0529. Such low values correspond to discs around black holes with masses of , whose SEDs visibly turn down towards the far-UV part of the spectrum. For example, Lai et al. (2023) show the red UV spectrum of their object J2157 that is in stark contrast to the blue spectrum of J0529, which implies a larger value for J0529 and suggests a lower-mass black hole for common stationary disc models. Also, using the mean relation from Netzer (2019), the luminosity of J0529 puts it far outside of the parameter space covered by the physically allowed range of black hole spin, while extrapolating their model calculation for any allowed spin to the correct luminosity puts it not only into a clearly super-Eddington regime but also suggests .
[…]
5.4 Predictions of UV line profiles
Using our BLR model we can predict the spectral line profile of the O I and C IV lines that were presented in Sec. 3.2. We have done this in a simplified way because our model is based on dynamics rather than photoionisation, hence we infer the location at which the corresponding line emission is produced by adjusting the size of the BLR for each line based on the ratio of the time lag of that line with respect to H as measured empirically from RM campaigns. For O I the scaling is because, as noted in Sec. 3.2, it is produced by Ly fluorescence and so is expected to arise at a similar location to the H I lines. Other low excitation lines such as Mg II and Fe II also have broadly similar time lags to H (Maoz et al., 1993; Hu et al., 2015; Prince et al., 2023; Shen et al., 2024). For simplicity we also assume that the O I line originates in the disky part of the BLR. For C IV, we adopt a time lag that is 30% of that for H, although there is again a large range of variation and the ratio may depend on luminosity (Lira et al., 2018; Shen et al., 2024). In addition the shape of the cross-correlation function may differ significantly between the lines (Horne et al., 2021). We assume C IV arises only in the outflowing part of our model. The resulting line profiles in Fig. 7 are similar to what one might expect given the contributions of disky and outflowing regions to the H line in Fig. 5. However, an important difference is the stronger and more extended blue wing on the predicted C IV profile, which qualitatively matches the observation shown in the lower panel of Fig. 3, and is a result of the way that decreases with radius. The higher excitation C IV originates from smaller radii and, lending support to the decreasing velocity profile we have adopted, has a higher outflow velocity.
@@ 5.5 Outflow rate
Following Shimizu et al. (2019), we estimate the outflow rate from our model as
@@
where is the radial range over which the estimate is made. We have applied this in a simplified way to calculate a single outflow rate between and the mean radius at 0.9 pc. For we use the mean radial velocity of the clouds within this range, and we assume a canonical density of cm-3. The resulting outflow rate, based on the inferred H luminosity, is M⊙ yr-1. We can apply a factor 2 correction to compensate the anisotropy in the model and another factor 2 because of the mid-plane opacity to account for the reverse outflow, increasing the implied outflow rate to M⊙ yr-1. This is very small compared to the accretion rate of M⊙ yr-1 required to maintain the luminosity, assuming a radiative efficiency of for super-Eddington sources (Jiang et al., 2019). However, it represents only the outflow of ionised gas in the BLR, and does not include the accretion disk wind launched on smaller scales, nor the radiation driven dusty wind from larger scales (which has the additional impact of reducing the inflowing mass that reaches the accretion disk). The lower outflow rate associated with the BLR is consistent with what was found by Vietri et al. (2018) for highly luminous QSOs. These authors showed that the kinetic power of the outflow from C IV in the BLR was an order of magnitude or more lower than that from [O III] on kiloparsec scales; and that winds from the accretion disk dominate the outflow on small scales to make up the difference.
@@ 6 Conclusions
We have used the VLTI interferometer GRAVITY+ to spatially resolve the kinematics of the BLR of the luminous z=4 QSO SMSS J052915.80-435152.0. We measured the differential phase spectra across both the H and H lines and combined these with a detailed flux spectrum from ERIS. We applied an updated modelling procedure that has a disky component and a separate outflowing component; and we fit both the flux profiles and phase spectra of the two H I lines simultaneously, allowing only the radial profiles of their line emission distribution to vary. The systemic velocity in the model differs by km s-1 from peak of the broad line emission, implying a redshift for the QSO of . Based on the model, we conclude:
@@
The ratio (of the strength of the iron to H emission) would imply a rather modest accretion rate, contrary to other evidence indicating there is a high accretion rate. We speculate that this discrepancy might be attributable to the dominance of the outflow in the BLR. If the scaling relations are applied using only the disky part of the BLR profile, the BH mass inferred from scaling relations is reduced, and the implied accretion rate increases. However, by comparing Fig. 2 and the bottom panels of Fig. 5 we caution that the decomposition of the BLR profile into outflowing and disky parts is dependent on geometry and anisotropy parameters, so cannot be done in a naive way by fitting a pair of Gaussian profiles.
An estimate of the outflow rate of ionised gas in the BLR is M⊙ yr-1. This is only a few percent of the accretion rate, indicating it is only a minor part of the total outflow rate which includes winds launched on both smaller and larger scales.

### v2

@@ 5.2 BLR geometry and size
The fitted parameters in Table 4 show that % of the clouds are in the outflow, while the remaining 17% are in the disky component on bound elliptical orbits. This disparity could indicate that the outflow dominates the BLR. But an alternative interpretation of this effect could be that the outflow is obscuring the disky part of the BLR. The model cannot distinguish between these options: it fits the properties of the observable line emission but does not address the physical processes that generate it.
@@
Our line of sight is near the top edge of the outflow, because the inclination is from face-on while the total angle from the mid-plane subtended by BLR clouds is . If the line of sight were oriented further into the outflowing wedge, one might expect to see broad absorption lines in the rest-frame optical spectrum (elv00) but these are not observed (wol24). The emission is anisotropic with indicating that the side of the illuminated side of each cloud emits preferentially. Equally important is the mid-plane opacity with meaning that the emission from clouds behind the mid-plane is almost entirely blocked, consistent with the basis of the FRADO model in which a dusty disk extends to small radii at low scale heights, and with the optically thick nature of super-Eddington accretion disks. This leads to the configuration in Fig. 5, where the clouds are coloured according to their line-of-sight velocity and have a size corresponding to their apparent brightness.
@@
The average radius for the BLR is pc for H and pc for H. While these are the same within their uncertainties (see Table 4 and Fig. 8), this difference is expected both observationally, based on RM delays for different lines (ben10), and on modelling of radiation pressure confined BLR clouds (net20).
@@
While the size appears to match that inferred by applying a correction to the R-L relation favoured by du19, this does not provide an explanation of why is so low. Indeed, from the plot of FWHM versus in she14, both of these values for J0529 seem rather typical of QSOs generally. Instead we propose that much of the deviation from the scaling relations that has previously been attributed to luminosity or accretion rate, could, analogous to the C IV line, be due to the impact of the outflow on the H I line properties. Indeed, the disky part of the BLR in Fig. 4 contains only 27% of the total line luminosity. Applying this to the typical equation which has a typical form where is usually taken to represent the total AGN luminosity. If instead we consider only the luminosity associated with the disky part of the BLR tracing rotationally supported gas (using the line luminosity as a suitable proxy), then the inferred size is reduced by a factor . This would bring the estimate from ben13 into good agreement with our measurement.
@@
Similarly, we would find which is what one would expect for a highly accreting QSO, particularly in the context of Eddington ratio discussed below. However, applying this new value would lead to a revised BLR size from du19 that is much too small. We therefore speculate that the over-estimate of the BLR size that has been discussed in the literature for highly accreting and highly luminous QSOs could be related to the impact of BLR outflows on the line properties.
@@ 5.3 Dynamical black hole mass
The form of the scaling relation used to estimate black hole mass is typically where represents AGN luminosity and (or FWHM) the line width. If, as above, we take to be a proxy for the line luminosity (noting that line luminosity is used explicitly in some relations), applying this to only the disky part of the BLR profile which has km -1 and contains 27% of the luminosity, reduces the inferred by a factor 14 (or a factor 5 if using FWHM, due to the differing FWHM/ ratios). This use of the rotational velocity of the BLR rather than the full line width (which is broadened by the outflow) would bring all the single epoch estimates much closer to the BH mass we have measured, and supports the speculation above that the outflow could be a major contributor to the bias in single epoch estimates for luminous QSOs.
@@ 5.4 Eddington ratio
The Eddington ratio is determined by the black-hole mass and the bolometric luminosity , which is commonly derived from a monochromatic luminosity and a bolometric correction . Also, the emission from the accretion disc requires an anisotropy correction that depends on the viewing angle, and to a smaller extent on the black-hole spin since the deflection of light due to gravitational lensing is stronger for disc annuli that are closer to the black hole (nem10; lai23). Generic values based on a single average spectral energy distribution (SED; ric06) are usually appropriate when the disc is large compared to the innermost stable circular orbit (ISCO) around the black hole. While the overall size and surface area of a disc is set by its luminosity, for black holes of increasing mass the hole inside the ISCO grows proportionally to mass. Discs around higher-mass black holes thus lack the hottest and most UV-emitting parts, which are present in discs of similar luminosity around lower-mass black holes. This implies a non-linear dependence of on mass and disc luminosity, which was explored in the framework of thin-disc models by several authors (e.g. nem10; net19; lai23). These explorations tend to offer a mean relation for as a function of monochromatic luminosity assuming a mean related black-hole mass or Eddington ratio.
@@
For J0529, we use the 5100 Å luminosity from Sec. 3 and the 33∘ inclination implied by our model. The mean relations by the three aforementioned works result in values ranging from 2.2 to 2.7, but are not applicable in the case of J0529. Such low values correspond to discs around black holes with masses of , whose SEDs visibly turn down towards the far-UV part of the spectrum. For example, lai23 show the red UV spectrum of their object J2157 that is in stark contrast to the blue spectrum of J0529, which implies a larger value for J0529 and suggests a lower-mass black hole for common stationary disc models. Also, using the mean relation from net19, the luminosity of J0529 puts it far outside of the parameter space covered by the physically allowed range of black hole spin, while extrapolating their model calculation for any allowed spin to the correct luminosity puts it not only into a clearly super-Eddington regime but also suggests .
Given the observed SED of J0529, is clearly too low, while we also believe that the canonical average-SED case of from ric06 is presumably too high. This point cannot be settled without data on the intrinsic disc emission at nm. Using these corrections, we infer conservatively that and hence which is substantially super-Eddington. In this regime, the luminosity increases only logarithmically with mass accretion rate (abr88), and there have been suggestions that the luminosity may saturate at a few to ten times due to the decreasing radiative efficiency (wat06; sad09). However, it might be possible to reach higher Eddington ratios: efficiencies that were unexpectedly high, exceeding those predicted by slim disc models, were reported by sad15 both for non-spinning, and even more so for spinning, black holes. In the latter case, the authors suggested that rotational energy of the black hole was extracted by a process similar to the Blandford-Znajek mechanism, giving rise to the question of whether there would be an associated jet. The only radio measurements available for J0529 is from the Rapid ASKAP Continuum Survey (RACS; mcc20), being 0.8 mJy at 0.9 GHz and 0.6 mJy at 1.4 GHz (D. McConnell, priv. comm.). Adopting the radio loudness parameter of kel89 as the ratio of the 6 cm flux density to that at 4400 Å (both measured in mJy), we find an intermediate value of (which barely changes if one adjusts the rest-frame wavelengths to account for the typical redshift of of the kel89 sample). This transitional value suggests that J0529 has a jet, but also implies that the optical continuum remains dominated by thermal disk emission.
ina20 noted that the accuracy of the numerical algorithms plays an important role in the outcome of the simulations, with reference to calculations by jia14; jia19. In particular, jia19 state that at accretion rates of the radiative efficiency remains high enough at % that one can have radiative luminosities of , but once the accretion rate reaches the radiative efficiency drops to below 1%. The super-Eddington accretion also leads to a thick disk with large scale heights (e.g. , jia19; jia25; hop25a) consistent with what we find in our model. Thus, J0529 may be characteristic of what is needed to grow massive black holes in the early universe (ina20): an extreme but physically plausible case of a QSO emitting close to its radiative limit in a regime where simulations indicate strong outflows are expected in a polar conical region.
@@ 5.5 Outflow rate
Following shi19, we estimate the outflow rate from our model as
@@
where is the radial range over which the estimate is made. We have applied this in a simplified way to calculate a single outflow rate between and the mean radius at 0.9 pc. For we use the mean radial velocity of the clouds within this range, and we assume a canonical density of cm-3. The resulting outflow rate, based on the inferred H luminosity, is M⊙ yr-1. We can apply a factor 2 correction to compensate the anisotropy in the model and another factor 2 because of the mid-plane opacity to account for the reverse outflow, increasing the implied outflow rate to M⊙ yr-1. This is very small compared to the accretion rate of M⊙ yr-1 required to maintain the luminosity, assuming a radiative efficiency of for super-Eddington sources (jia19). However, it represents only the outflow of ionised gas in the BLR, and does not include the accretion disk wind launched on smaller scales, nor the radiation driven dusty wind from larger scales (which has the additional impact of reducing the inflowing mass that reaches the accretion disk). The lower outflow rate associated with the BLR is consistent with what was found by vie18 for highly luminous QSOs. These authors showed that the kinetic power of the outflow from C IV in the BLR was an order of magnitude lower than that from [O III] on kiloparsec scales; and that winds from the accretion disk dominate the outflow on small scales to make up the difference.
6 Comparison to the UV lines C IV and O I
Differences between the high ionisation lines (HILs) and low ionisation lines (LILs) in the spectra of QSOs have been reported and analysed extensively in the literature for more than four decades. In one of the first papers to address this topic, gas82 showed that the HILs were blue-shifted by typically 600 km s-1 with respect to the LILs, due to radial motions. To explain this phenomenon, a physical model was proposed by col88 in which the LIL emission comes from clouds in the outer part of the accretion disk, while the HIL emission originates in outflowing transient clouds. A better understanding of the physical situation came with higher quality spectra: wider band coverage, higher resolution, and better signal-to-noise. In particular, detailed UV spectra of two local Narrow-Line Seyfert 1 AGN were presented by lei04a. These highlighted that the HILs were both strongly blue-shifted and asymmetric, while the LILs were narrow, symmetric, and at the systemic velocity of the AGN. An extensive effort building on the eigenvector classification for local QSOs (bor92) led to incorporation of the C IV line into the scheme, bringing it to 4-dimensions (sul11; sul15; sul17). It is now generally accepted that the line properties can be grouped into two classes of QSOs related to whether the C IV is dominated by a more systemic or strongly blue-shifted component. (mar17; vie18; vie20; dec23).
Despite this, C IV has been used to estimate because it is easily observable in high redshift QSOs and is known to originate in the BLR (ves06; par13; run13). While this is physically reasonable for QSOs in which the line is both approximately symmetric and close to the systemic velocity, it is harder to explain a relation between the line properties and for emission originating in an outflow. Nevertheless, coa17 found an empirical relation between the line’s blue-shift and a correction factor that can be applied to the inferred mass, to match it to other scaling relations. However, tem23 have shown that the situation is more complex, and that there is a 2-dimensional parameter space relating the blue-shift (and equivalent width) to a combination of and . They showed there is a wedge in that parameter space where the C IV blue-shift increases (and equivalent width decreases), and that in general the highest blue-shifts are observed in AGN with both and . Their modelling to reproduce these effects is bringing us to a better understanding of the cause of the two populations of QSOs with differing line properties noted above.
In Fig. 6 we compare the H I line profile of J0529 to the C IV line, as well as the O I in the Xshooter spectrum published by wol24. To do so, we have subtracted an estimate of the true continuum level as indicated by the lowest regions in the spectral ranges around the two lines. We have not tried to fit the iron emission, nor other neighbouring emission, since the aim is only to compare the line profiles.
6.1 Observation and Model of C IV
The C IV 1549 Å line in the upper left panel of Fig. 6 appears to be mostly associated with the outflow, and can be traced to a velocity of . It has a blue-shift of km -1, which is higher than the sample range of tem23 but consistent with their analysis based on the high and , and confirms the nuclear nature of the outflow.
Using our BLR model we can make a simple prediction of the spectral profile of the line. Because our model is based on dynamics rather than photoionisation, we can do this only in a crude way by adjusting the size of the BLR emission for the line according to an estimate of its expected radial location from RM measurements. From the size-luminosity relation of the C IV line and 1350 Å continuum (based on 7-13 yr RM campaigns of QSO samples) we estimate a characteristic radius for the C IV emission IN J0529 of 0.3 pc (lir18; kas21; she24). We note, however, that there is large range of variation between relations including a possible dependence on luminosity; and in addition the shape of the cross-correlation function may differ significantly between the C IV and H I (hor21).
For this comparison, we have assumed that C IV arises only in the outflowing part of our model. Although this is physically unlikely, the reason for doing so is that in the BLR model, both the disky and outflowing components extend over the same radial range with the same radial profile. To modify the ratio of their contributions would require manually adjusting the scaling between them. Although doing so would create a better match to the overall profile, it would not lead to further insights, and so we have chosen to focus on a comparison of the outflowing part.
The resulting line profile is shown in the lower left panel of Fig. 6, where the stronger and more extended blue wing on the predicted C IV profile, which qualitatively matches the observation, is a result of the way that decreases with radius. The higher excitation C IV originates from smaller radii and, lending support to the decreasing velocity profile we have adopted, has a higher outflow velocity.
6.2 Observation and Model of O I
The upper right panel of Fig. 6 compares the H I line profile to O I 1302 Å, a low excitation line that in AGN is expected to arise in the BLR through Ly fluorescence (net79; kwa81; mat07). O I 8446 Å is also enhanced through this process and the ratio between these lines in the resulting radiative cascade can provide a measure of the extinction towards the BLR (net79). Given that the Balmer line ratios in Sec. 3 hint that there might be some modest extinction, an independent measure like this would enable one to clarify the impact. There is a good match on the red side of the profiles, but the O I line shows less evidence for a blue wing suggesting it is more likely associated with the systemic velocity of the QSO. We note that the peak of the O I line is contaminated by deep absorption across a range of velocities, the cause of which is unclear.
With the same limitations and cautions as above for C IV, we can attempt to reproduce the O I profile using our model of the BLR. Because the emission is produced by Ly fluorescence, it is expected to arise at a similar location to the H line. We have therefore set the radial scaling for the BLR to , noting that RM campaigns have shown that other low excitation lines such as Mg II and Fe II also have broadly similar time lags to H (mao93; hu15; pri23; she24). For simplicity we also assume that the O I line originates only in the disky part of the BLR, although we acknowledge there is some indication of a weak outflow component in the observed profile. The predicted O I line is shown in the lower right panel of Fig. 6. In particular, the match of the red side of the profile tends to confirm that the emission is dominated by the disky part of the BLR; and, due to the difference of the blue side of the profile, that some emission must arise in the outflow.
@@ 7 Conclusions
We have used the VLTI interferometer GRAVITY+ to spatially resolve the kinematics of the BLR of the luminous z=4 QSO SMSS J052915.80-435152.0. We measured the differential phase spectra across both the H and H lines and combined these with a detailed flux spectrum from ERIS. We applied an updated modelling procedure that has a disky component and a separate outflowing component; and we fit both the flux profiles and phase spectra of the two H I lines simultaneously, allowing only the radial profiles of their line emission distribution to vary. While the redshift of the QSO inferred from the broad line peak is , the systemic velocity for the best fit differs from that by km s-1 and favours a redshift of within the scope of the model. Based on the model, we conclude:
@@
The ratio (strength of the iron to H emission) would imply a rather modest accretion rate, contrary to other evidence indicating there is a high accretion rate. We speculate that this discrepancy might be attributable to the dominance of the outflow in the BLR. If the scaling relations are applied using only the disky part of the BLR profile, the BH mass inferred from scaling relations is reduced, and the implied accretion rate increases. However, by comparing Fig. 2 and the bottom panels of Fig. 4 we caution that the decomposition of the BLR profile into outflowing and disky parts is dependent on geometry and anisotropy parameters, so cannot be done in a naive way by fitting a pair of Gaussian profiles.
An estimate of the outflow rate of ionised gas in the BLR is M⊙ yr-1. This is about 1% of the accretion rate, indicating it is only a minor part of the total outflow rate that includes winds launched on both smaller and larger scales.